OpenStudy (anonymous):

What is the third term of (a + b)^11?

OpenStudy (campbell_st):

if you know about binomial expansion using combination notation it is \[^{n}C_{r} a^{n - r}b^{r}\] in your question n = 11 r = 2

OpenStudy (anonymous):

can you explain idk what is binomial expansion

OpenStudy (hba):

Some good notes would help.

OpenStudy (campbell_st):

ok... do you know about pascals triangle...?

OpenStudy (campbell_st):

well that makes it easy... go and read the notes... its impossible to do a problem if you don't have an understanding of the key concepts...

OpenStudy (anonymous):

yea i do

OpenStudy (anonymous):

is this correct ? C(11,2) a^9 b^2

OpenStudy (campbell_st):

it is correct... you just need to evaluate the combination \[^{11}C_{2} = \frac{n!}{r!(n - r)!}= \frac{11!}{2!(11 - 2)!}\]

OpenStudy (anonymous):

so i need to solve 11!/2!(11 - 2)!? right